Number of line per centimeter of a plan diffraction grating is 6000.when a monochromatic light is incident on it the angle of diffraction for the second order spectral line become 450.find the wavelength of the light.


1
Expert's answer
2020-12-23T07:46:08-0500

Answer

equation for diffraction grating

"d\sin\theta=n\lambda"

No. Of spectral line per centimeter

1/d=6000

In first case

"\lambda=\sin\theta/6000"

Now in second case

Wavelenth

"\lambda'=\frac{\sin\theta}{6000\times2\times sin\theta\times450}=0.9A^°"